We want to form a license plate using 4 letters (A-Z) followed by 3 numbers (0-9). Both may be used more then once. A license plate is chosen randomly. What is the probability that the letter A is not used?
Answer by stanbon(75887)

We want to form a license plate using 4 letters (A-Z) followed by 3 numbers (0-9). Both may be used more then once. A license plate is chosen randomly. What is the probability that the letter A is not used?
----
Number of plates with no "A" = 25^4*10^3
----
Number of plates with no restriction = 26^4*10^3
-----------------------------------------------------
P(letter A not used) = 25^4/26^4 = 0.8548
========================
Cheers,
Stan H.